# Toward Efficient and Robust Behavior Models for Multi-Agent Driving Simulation

### v1

@@ III-D Behavior Model Details
where $\zeta(\cdot)$ and $\beta(\cdot)$ are both realized as MLP s. This operation yields the representation of token $\mathbf{z}_{j}$ relative to instance ii.
@@ IV Experiments
Datasets: We evaluate our approach on two datasets: 1) the publicly available INTERACTION dataset [zhan2019interaction], containing 38 25538\,255 trajectories ($4.15\text{\,}\%$ VRUs) across $11$ locations; and 2) an in-house dataset commercially licensed by DeepScenario, comprising 88 54088\,540 trajectories ($43.59\text{\,}\%$ VRUs) across 88 locations. Both datasets were captured using drones, ensuring high-quality data, and include highly interactive scenarios such as merging lanes, unsignalized intersections, and roundabouts. We use $20\text{\,}\%$ of the recordings per location for validation and $30\text{\,}\%$ for testing. A more compact version of the DeepScenario dataset [dhaouadi2025deepscenario] is publicly available and can be used to obtain similar results.
@@
For our proposed model, we use 33 Perceiver layers with a hidden dimension of $128$. We also evaluate a more compact variant with a single Perceiver layer and hidden dimension $64$, denoted Ours (small). All baseline models use a hidden dimension of $128$.
@@ IV-A Results
Effects of Average Reward Target: Figure 7 presents the results of our ablation study on different choices for the average reward target $\tilde{r}_{\mathrm{target}}$, evaluated across values from 11 to $29$ in increments of 22. In general, increasing $\tilde{r}_{\mathrm{target}}$ leads to more robust behavior, with larger gains being achieved at lower values. Initially, the positional accuracy, measured by RMSE, increases significantly, demonstrating that the model benefits from the improved robustness. However, increasing $\tilde{r}_{\mathrm{target}}$ too much, leads to the positional error to rise again. For instance, this turning point can easily be seen for $\tilde{r}_{\mathrm{target}}=11$ on the DeepScenario dataset. This indicates a shift in the model’s behavior, where it starts to prioritize robustness over realism. Based on these results, we selected the values $11$ for DeepScenario and 1919 for INTERACTION, representing the optimal trade-off between prediction accuracy and robustness.

### v2

@@ III-D Behavior Model Details
where $\zeta(\cdot)$ and $\beta(\cdot)$ are both realized as MLP s. This operation yields the representation of token $\mathbf{z}_{j}$ relative to instance $i$.
@@ IV Experiments
Datasets: We evaluate our approach on two datasets: 1) the publicly available INTERACTION dataset [zhan2019interaction], containing 38 25538\,255 trajectories ($4.15\text{\,}\%$ VRUs) across $11$ locations; and 2) an in-house dataset commercially licensed by DeepScenario, comprising 88 54088\,540 trajectories ($43.59\text{\,}\%$ VRUs) across $8$ locations. Both datasets were captured using drones, ensuring high-quality data, and include highly interactive scenarios such as merging lanes, unsignalized intersections, and roundabouts. We use $20\text{\,}\%$ of the recordings per location for validation and $30\text{\,}\%$ for testing. A more compact version of the DeepScenario dataset [dhaouadi2025deepscenario] is publicly available and can be used to obtain similar results.
@@
For our proposed model, we use $3$ Perceiver layers with a hidden dimension of $128$. We also evaluate a more compact variant with a single Perceiver layer and hidden dimension $64$, denoted Ours (small). All baseline models use a hidden dimension of $128$.
@@ IV-A Results
Effects of Average Reward Target: Figure 7 presents the results of our ablation study on different choices for the average reward target $\tilde{r}_{\mathrm{target}}$, evaluated across values from $1$ to $29$ in increments of $2$. In general, increasing $\tilde{r}_{\mathrm{target}}$ leads to more robust behavior, with larger gains being achieved at lower values. Initially, the positional accuracy, measured by RMSE, increases significantly, demonstrating that the model benefits from the improved robustness. However, increasing $\tilde{r}_{\mathrm{target}}$ too much, leads to the positional error to rise again. For instance, this turning point can easily be seen for $\tilde{r}_{\mathrm{target}}=11$ on the DeepScenario dataset. This indicates a shift in the model’s behavior, where it starts to prioritize robustness over realism. Based on these results, we selected the values $11$ for DeepScenario and $19$ for INTERACTION, representing the optimal trade-off between prediction accuracy and robustness.